# Why Fine-Tuning Struggles with Forgetting in Machine Unlearning? Theoretical Insights and a Remedial Approach

### v1

@@ B.2 Proof of Theorem 1
Let us first focus on the performance of the golden model. Based on the definition of unlearning accuracy and remaining accuracy, we have
@@
=$\overset{(i)}{=}\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{P}\mathbf{w}_{*}^{r}+\mathbf{X}_{f}^{\top}\mathbf{P}\mathbf{w}_{*}^{f}-\mathbf{X}_{f}^{\top}\mathbf{w}_{*}^{f}]\|^{2}$
=$\overset{(ii)}{=}0,$
where $(i)$ follows from $\mathbf{X}_{f}^{\top}\mathbf{P}_{t}=\mathbf{X}_{f}^{\top}\mathbf{P}_{r}=0$, and $(ii)$ follows from $\mathbf{X}_{f}^{\top}\mathbf{P}=\mathbf{X}_{f}^{\top}$. The proof is finished.
@@ B.3 Proof of Theorem 2
For the fine-tuning, the remaining accuracy follows:
@@
$=0,$
where the last equality follows from: (i) the projection matrix is a commutative matrix; (ii) the last property in Corollary 3. For the unlearning accuracy, it holds that
@@
=$\overset{(i)}{=}\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{P}_{t}\mathbf{w}_{*}^{f}]\|^{2}$
=$\overset{(ii)}{=}0,$
where $(i)$ follows from Corollary 3, and $(ii)$ follows from the fact that $\mathbf{X}_{t}$ has the same data structure as $\mathbf{X}_{t}$ such that:
@@ B.4 Proof of Theorem 3
For the non-overlapping case, we have that the remaining accuracy follows:
@@
$=\frac{1}{n_{f}}\|\mathbf{w}_{*}^{f}]\|_{\mathbf{X}_{f}\mathbf{X}_{f}^{\top}}^{2}.$
@@
Then, the proof is complete.

### v2

@@ B.2 Proof of Theorem 1
Let us first focus on the performance of the golden model. Based on the definition of unlearning accuracy and retaining accuracy, we have
@@
$=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{P}\mathbf{w}_{*}^{r}+\mathbf{X}_{f}^{\top}\mathbf{P}\mathbf{w}_{*}^{f}-\mathbf{X}_{f}^{\top}\mathbf{w}_{*}^{f}]\|^{2}$
$=0,$
where the penultimate equality comes from $\mathbf{X}_{f}^{\top}\mathbf{P}_{t}=\mathbf{X}_{f}^{\top}\mathbf{P}_{r}=0$, and the last equality follows from $\mathbf{X}_{f}^{\top}\mathbf{P}=\mathbf{X}_{f}^{\top}$.
@@ B.3 Proof of Theorem 2
For the fine-tuning, the retaining accuracy follows:
@@
$=0.$
The last equality derives from that the facts the projection matrix is commutative matrix and the last property holds in Corollary 3. For the unlearning accuracy, it holds that
@@
$=\frac{1}{n_{f}}\|\mathbf{X}_{f}^{\top}\mathbf{P}_{t}\mathbf{w}_{*}^{f}]\|^{2}$
$=0,$
where the penultimate equality is due to Corollary 3 and the last equality comes from the fact $\mathbf{X}_{t}$ enjoys the same data structure as $\mathbf{X}_{t}$ such that:
@@ B.4 Proof of Theorem 3
For the non-overlapping case, we have that the retaining accuracy follows:
@@
$=\frac{1}{n_{f}}\|\mathbf{w}_{*}^{f}]\|_{\mathbf{X}_{f}\mathbf{X}_{f}^{\top}}^{2},$
@@
The proof is then complete.